1. If the atomic radius of sodium atoms is 1.86 × 10−10 m, how many sodium atoms are needed to make a line that is 1.00 cm in length?
1
Expert's answer
2021-09-23T02:37:12-0400

= 1.86 × 10-10×100 = 1.86 × 10-8

= 1/1.86 × 10-8 = 5.376344086 × 107 sodium atoms